HOPE operational status:
The station is back up after the station operator replaced the Newmar power supply.
The station is down due to a damaged Newmar power supply. IDA engineers have prepared replacements to be shipped.
The station operator replaced the damaged uplink cable on day 2022,061.
The station was down from day 2021,216 to 218 due to a short in the uplink cable between the vault and recording room.
The station operator repaired the uplink cable and restored data acquisition at the station. There is further damage to the cable and it will need to be replaced.
The station is down as of day 2021,126 (May 6) due a short in the uplink cable causing loss of communication between the vault and recording room.
Telemetry at the station is back up and data acquisition was restored on day 2017,048. There is a data gap from day 2017,017-14:12 to 2017,048-12:17.
Telemetry is down since 2017,010-11:29 due to damaged VSAT equipment. The circuit will be down until replacements can be shipped.
IDA station HOPE was down from day 2015,008-16:55 to 2015,017-11:43. An IDA engineer is on-site upgrading the vault. Additional work includes installing a Setra microbarograph, Episensor strong motion sensor, and upgrading the power equipment.
The station is back on-line after an IDA engineer visited the station and upgraded it to a Q330HR system. Seismometers include an STS-2, Trillium, and Episensor strong motion sensor.
Data quality for all channels from the STS-2 is poor since day 2010,055-02:30 for reasons yet to be determined. The channels are showing excessive spiking and noise. A trip to upgrade the station is planned, pending shipment of equipment to the station.
The station operator re-centered the STS-2, restoring the channels at 2010,311-15:33.
The station is down as of day 2010,306-03:22. All channels from the STS-2 and FBA-23 sensors are non-seismic for reasons yet to be determined.
The station was down from day 2010,216-16:49 to 2010,219-12:48 due to problems with the DAS.
All channels from the STS-2 seismomter are showing large spikes every few hours starting at about day 2010,068-08:10, for reasons yet to be determined. The sensor will be replaced by an IDA engineer during the next station visit.
The STS-2 sensors have been reconfigured to record channels in high gain (24 bit) resolution. Channel names and sample rates have been updated accordingly.
The STS-1 sensors have been removed from the station as of about day 2007,296.
The East-West channels from the STS-1 sensor are seismic again as of about day 135-10:00. However, the mass position channels are still flatlined.
The East-West (BBE) channels from the STS-1 sensor have drifted back out of normal operating range as of about day 133-03:26.
The East-West (BBE) channels from the STS-1 sensor have drifted back into normal operating range about day 127-20:05. However, the North-South channels for the STS-1 have drifted out of range, and the mass position channel is flatlined as of about day 128-18:10.
The North-South channels from the STS-1 sensor drifted back into normal operating range about day 031-01:43. However, the masses for all STS-1 channels remain in an unstable state and will need to be recentered.
Data acquisition resumed about day 344-15:21 after the station operator cycled power to the ISP, and an IDA engineer rebuilt the disk loops.
IDA station HOPE is down as of about day 342-10:53 due to communication problems between the DAS and the ISP that are yet to be determined. The vertical channels from the STS-1 are non-seimic as of 340-18:00. The mass position for all the STS-1 sensors need to be recentered.
The E-W channels for the STS-1 sensor have drifted out of normal operating range at about 234-08:45. The mass position channels for the STS-1 sensors are in an unstable state and need to be recentered.
The N-S channels for the STS-1 sensor have drifted out of normal operating range and the masses cannot be recentered due to a faulty motor in the sensor.
An IDA engineer visited the station and upgraded the station telemetry to a VSAT system, replaced the ISP, and adjusted the masses for the STS1 sensors.
Station operator reports DAT drive failure. The station operator recentered the masses for the STS1 sensors. The E-W is back to normal operating range. However, the N-S still needs adjustment.
Masses for both horizontal STS1 sensors seem in an unstable state. The free period on all three STS1’s masses needs to be adjusted by an IDA engineer on the next station visit.
The masses of the STS1 sensors have been adjusted back to their normal range.
The masses of the STS1 sensors (N-S and vertical) have drifted out of the operating range and will have to be manually readjusted.
The North-South channels (BBN), and vertical channel (VMZ), for the STS1 sensor are out following a reboot of the DAS. Engineering tests conducted 2003.07.18 suggest a problem with the sensor.
Connection to station restored.
Lost connection to station.
New EPROMs installed to correct DSP problem. New NRTS workstation installed.
Station is operating normally. Data acquisition restored on 2002.04.15 during visit by field engineer. Uplink cable had been severed during relocation of host institution facilities.